Micheal drove to a friends house at a rate of 40 mi/h. He returned by the same route at a rate of 45 mi/h. The driving time for the round trip was 4 h. What is the distance Michael traveled.
Answer by Alan3354(69443)

Micheal drove to a friends house at a rate of 40 mi/h. He returned by the same route at a rate of 45 mi/h. The driving time for the round trip was 4 h. What is the distance Michael traveled.
------------
Avg speed for the round trip = 2*40*45/(40+45) = 3600/85 mi/hr
Avg speed = 720/17 mi/hr
-----
d = r*t
d = (720/17)*4 = 2880/17 miles round trip